# Solute Transport due to Periodic Loading in a Soft Porous Material

### v1

@@ III.6 Péclet number in biological examples
We conclude by reporting typical loading parameters, material properties and diffusion coefficients for several examples of soft biological tissues (table III.6). Based on these values, we calculate the Péclet number as defined in section II.7, which quantifies the importance of poromechanical relaxation relative to molecular diffusion, and an effective Péclet number $\mathrm{Pe_{eff}}=\mathrm{Pe}A/T$, which additionally accounts for the impact of loading on advection to understand the range of loading amplitudes and frequencies where advection and dispersion are likely to be non-negligible. For simplicity, we use a moderate dimensionless amplitude $A=0.1$ in all cases. In all cases, the values of $\mathrm{Pe_{eff}}$ range from moderate to high, indicating that advection and potentially dispersion are probably not negligible in these systems.
[…]
Brain ECM Kedarasetti et al. (2020) $2\times 10^{-4}$ $0.1$ $2\times 10^{-15}$ $2\times 10^{3}$ $10$ 0.30.30.30.310101010 $0.003\pi$$0.1\pi$ $1.4\times 10^{-10}$ $30$ $3\times 10^{1}$$10^{3}$
Cartilage Ferguson et al. (2004); Gardiner et al. (2007); DiDomenico et al. (2017) $2\times 10^{-3}$ $0.1$ $7.5\times 10^{-18}$ $5\times 10^{6}$ $10^{2}$ 0.0010.0010.0010.0010.10.10.10.1 (sitting) 0.10.10.10.11111 (running) $0.03\pi$$3\pi$ $0.003\pi$$0.03\pi$ $5\times 10^{-12}$ $7\times 10^{3}$ $10^{2}$$10^{3}$ $10^{3}$$10^{5}$
Intervertebral disk (Anulus F.) Ferguson et al. (2004) $10^{-2}$ $0.1$ $7.5\times 10^{-19}$ $2.5\times 10^{6}$ $5\times 10^{4}$ $2\times 10^{-5}$ (wake cycle) 0.0010.0010.0010.0010.10.10.10.1 (sitting) 0.10.10.10.11111 (running) $0.3\pi$ $6\times 10^{-5}\pi$$6\times 10^{-3}\pi$ $6\times 10^{-6}\pi$$6\times 10^{-5}\pi$ $5\times 10^{-12}$ $4\times 10^{2}$ $4\times 10^{1}$$2\times 10^{3}$ $2\times 10^{3}$$2\times 10^{5}$ $2\times 10^{5}$$2\times 10^{6}$
Cartilage scaffold (bioreactor) Sengers et al. (2004) $2\times 10^{-3}$ $0.1$ $10^{-17}$ $10^{5}$ $4\times 10^{3}$ 0.0010.0010.0010.0011111 $0.0001\pi$$0.1\pi$ $10^{-9}$ $10^{4}$ $4\times 10^{1}$$4\times 10^{4}$
@@ IV Conclusions
Our analysis is linked to a companion study, where we characterise the mechanical response of a soft porous medium under the same loading scenarios Fiori et al. (2023). That study showed that, depending on the loading period, the deformation can belong to either a slow-loading regime – in which the deformation is uniform throughout the domain – or a fast-loading regime – in which the deformation is increasingly localised near the permeable boundary (i.e., the left boundary). Fiori et al. (2023). Here, we analysed how these different mechanical behaviours related to the loading parameters influence solute transport and mixing.
@@
We then introduced two variables — travel distance $\delta$ and degree of mixing $\chi$ — and compared them for different transport coefficients and different loading parameters over several loading cycles. We showed that diffusion and dispersion are roughly Fickian on average. For advection and diffusion only (no dispersion), larger amplitudes increasingly reduce the average porosity and hence increasingly suppress diffusion. This case is essentially independent of $T$ except for very small values of $T$, where the material is additionally compressed and $\delta$ and $\chi$ decrease as $T$ decreases. When including dispersion, solute transport increases monotonically with $A$ but varies non-monotonically with $T$. The latter is due to the progressive localisation of the deformation at the left boundary, which reduces the intensity of the dispersive flux at the right boundary, where the solutes are positioned.
@@
We have taken the diffusion and dispersion coefficients themselves to be constants. As noted above, future work should consider the impact of the changing pore structure on the values of these material properties. It is likely that, much like $k$, both $D_{m}$ and $\alpha$ should be deformation dependent for moderate to large deformations.
@@ Appendix A Numerical method
Our model equations are solved using finite difference methods. We use a sixth-order compact finite difference approximation for spatial derivatives and a third-order compact finite difference approximation for derivative boundary conditions Lele (1992). For time integration, an implicit Runge-Kutta method is used via MATLAB’s built-in solver ODE15s Shampine and Reichelt (1997).
@@
$\frac{\partial\Phi}{\partial t}+\frac{\partial}{\partial x}[F(\Phi)]=0.$ (40)
@@
which maps $a\leq{}x\leq{}1$ to $10\leq{}\xi\leq{}$ 1. Equation 40 can then be rewritten in terms of $\xi$ as
@@
For our spatial discretisation, we perform a convergence analysis in the number of grid points $N_{x}$ (see figure 15) by calculating the Root Mean Square (RMS) relative eror in $c(t,x=1)$ for each solution with respect the solution for $N_{x}=1000$.
@@ Appendix B Analytical solution for molecular diffusion and advection
The analytical solution for the molecular diffusion of a delta function in a semi-infinite material, as formulated by Crank (1979), is:
@@
In case of slow-loading, the analytical solution for advection ($\mathrm{Pe}^{-1}=\alpha=0$) is
$\tilde{c}(\tilde{x},t)=\frac{1}{2}\bigg{\{}\tanh{\bigg{[}\tilde{s}\left((\tilde{x}-1)\frac{\phi_{f,0}}{\phi_{f}}+\tilde{l}\right)}\bigg{]}+1\bigg{\}}.$ (46)
@@ Appendix C Dispersive flux
In this Appendix, we justify the choice of fixing $\mathrm{Pe}^{-1}$ and $\alpha$ to the specific baseline values reported in III.2. We do so by quantifying the strength of dispersion relative to diffusion using the ratio of the dispersive solute flux $q_{\mathrm{disp}}$ to the diffusive solute flux $q_{\mathrm{diff}}$ which we define as
@@
Figure 17: (a) Lowest and (b) largest values of $q_{\mathrm{disp}}/q_{\mathrm{diff}}$ considered in this study. The flux is plot at ten different values of $X=x-u_{s}(X,t)$ from 00 to $1$ (dark to light) during one cycle. We distinguish between the loading half of the cycle ($\dot{a}>0$; solid curves) and the unloading half of the cycle ($\dot{a}<0$; dashed curves).
@@ Appendix D Initial porosity and solute strip width
Decreasing $\phi_{f,0}$ leads to larger oscillations in fluid flux for a given solid velocity (see eq 22) In this way, both advection and hydrodynamic dispersion are enhanced. However, the resulting impact on travel distance in figure 18 is relatively small because dispersion is relatively weak for these parameters.
Increasing the initial solute amount leads to a sharp increase in the travel distance: this is generally expected because the total concentration of solute is higher and therefore all the transport mechanisms are amplified. In addition, the flux fluid increases monotonically in magnitude from right to left, so larger values of l expose the solute front to stronger advection and dispersion. However, note that a wider solute strip is associated with a lower degree of mixing because the variance of the solute compared to the initial variance is lower when there is more solute (see equation 35).

### v2

@@ III.6 Péclet number in biological examples
We conclude by reporting typical dimensionless loading parameters and diffusion coefficients for several examples of soft biological tissues (table 1). Based on these values, we calculate the Péclet number and the effective Péclet number, as defined in section II.7, to understand the range of loading scenarios where advection and dispersion are likely to be non-negligible. For simplicity, we use a moderate dimensionless amplitude $\tilde{A}=0.1$ in all cases. The detailed parameters to calculate $\tilde{T}$ and $\mathrm{Pe}$ are reported, for the same biological tissues, in Fiori et al. (2023). In all cases, the values of $\mathrm{Pe_{eff}}$ range from moderate to high, indicating that advection and potentially dispersion are unlikely to be negligible in these systems.
Table 1: Loading and transport parameters for some examples of biological materials.
Tissue $\mathcal{D}_{m}$ [$\mathrm{m}^{2}\,\mathrm{s}^{-1}$] $\mathrm{Pe}$ $\tilde{T}/\pi$ $\mathrm{Pe_{eff}}=\mathrm{Pe}(\tilde{A}/\tilde{T})$
Brain ECM (Kedarasetti et al., 2022) $1.4\times 10^{-10}$ $30$ 0.0030.0030.0030.0030.10.10.10.1 $3\times 10^{1}$$1\times{}10^{3}$
Cartilage (Ferguson et al., 2004; Gardiner et al., 2007; DiDomenico et al., 2017a) $5\times 10^{-12}$ $7\times 10^{3}$ 0.030.030.030.033333 (sitting) 0.0030.0030.0030.0030.030.030.030.03 (running) $10^{2}$$10^{3}$ $10^{3}$$10^{5}$
Intervertebral disc (Annulus F.) (Ferguson et al., 2004) $5\times 10^{-12}$ $4\times 10^{2}$ $0.3$ (wake cycle) $6\times 10^{-5}$$6\times 10^{-3}$ (sitting) $6\times 10^{-6}$$6\times 10^{-5}$ (running) $4\times 10^{1}$$2\times 10^{3}$ $2\times 10^{3}$$2\times 10^{5}$ $2\times 10^{5}$$2\times 10^{6}$
Cartilage scaffold (bioreactor) (Sengers et al., 2004) $10^{-9}$ $10^{4}$ $10^{-4}$$10^{-1}$ $4\times 10^{1}$$4\times 10^{4}$
@@ IV Conclusions
Our analysis is linked to a companion study, where we characterise the mechanical response of a soft porous medium under the same loading scenarios Fiori et al. (2023). That study showed that, depending on the loading period, the deformation can belong to either a slow-loading regime – in which the deformation is uniform throughout the domain – or a fast-loading regime – in which the deformation is increasingly localised near the permeable boundary (i.e., the left boundary) Fiori et al. (2023). Here, we analysed how these different mechanical behaviours related to the loading parameters influence solute transport and mixing.
@@
We then considered two variables — travel distance $\delta$ and degree of mixing $\chi$ — and compared them for different transport coefficients and different loading parameters over several loading cycles. We showed that diffusion and dispersion are roughly Fickian on average. For advection and diffusion only (no dispersion), larger amplitudes increasingly reduce the average porosity and hence increasingly suppress diffusion. This case is essentially independent of $T$ except for very small values of $T$, where the material is additionally compressed and $\delta$ and $\chi$ decrease as $T$ decreases. When including dispersion, solute transport increases monotonically with $A$ but varies non-monotonically with $T$. The latter is due to the progressive localisation of the deformation at the left boundary, which reduces the intensity of the dispersive flux at the right boundary, where the solutes are positioned. We formalise these trends in the expression of the characteristic transport length $\delta_{est}$, which summarises the different regimes according to the transport and loading parameters.
For small deformations, it may be possible to derive an effective equation for net transport in this system. For example, Pool et al. (2016) used a multiple-scales technique to develop an effective transport equation for periodic loading in a semi-infinite domain that is valid for times much larger than the forcing period. The same approach could potentially be adapted to the present scenario, where the small-deformation analytical solution for the flow field is a full Fourier series rather than a single exponentially damped mode because the domain is bounded. This analysis is beyond the scope of the present work, but it would provide valuable qualitative insight into deformation-driven transport.
@@
We have taken the diffusion and dispersion coefficients themselves to be constants. As noted above, future work should consider the impact of the changing pore structure on the values of these material properties and, ultimately, on the underlying model for hydrodynamic dispersion. At the least, it is likely that, much like $k$, both $D_{m}$ and $\alpha$ should be deformation-dependent for moderate to large deformations.
Future work should also include the expansion of this study to 2D and 3D media with heterogeneous properties. Already for small deformations, periodic deformations have been shown to originate complex transport and mixing dynamics in heterogeneous media Pool and Dentz (2018); Trefry et al. (2019); Wu et al. (2020, 2024) and in response to multi-modal forcing (Trefry et al., 2020). Complex media and forcings are particularly relevant to biological systems; examples include heterogeneous structures originating in the multi-scale nature of tissues and/or multi-modal loading originating in the co-existence of different vital cyclic pulsations, such as the cardiac and the respiratory cycles. Finally, future work should also focus on an experimental validation of the results presented here.
@@ Appendix A Numerical method
Our model equations are solved using finite difference methods. We use a sixth-order compact finite difference approximation for spatial derivatives and a third-order compact finite difference approximation for derivative boundary conditions Lele (1992). For time integration, we use an implicit Runge-Kutta method via MATLAB’s built-in solver ODE15s Shampine and Reichelt (1997).
@@
$\frac{\partial\Phi(x,t)}{\partial t}+\frac{\partial}{\partial x}[F(\Phi(x,t)]=0.$ (44)
@@
which maps $a\leq{}x\leq{}1$ to $10\leq{}\xi\leq{}$ 1. Performing a standard change of variables from $(x,t)$ to $(\xi,t)$, partial derivatives transform according to
$\frac{\partial}{\partial{t}}\,\to\,\frac{\partial}{\partial{t}}+\frac{\partial{\xi}}{\partial{t}}\frac{\partial}{\partial\xi}\quad\mathrm{and}\quad\frac{\partial}{\partial{x}}\,\to\,\frac{\partial{\xi}}{\partial x}\frac{\partial}{\partial\xi}$ (46)
and equation (44) can be rewritten as
@@
For our spatial discretisation, we perform a convergence analysis in the number of grid points $N_{x}$ (see figure 15) by calculating the Root Mean Square (RMS) relative error in $c(t,x=1)$ for each solution with respect to the solution for $N_{x}=1000$.
@@ Appendix B Analytical solution for molecular diffusion and advection
The analytical solution for the molecular diffusion of a step function in a semi-infinite material, as formulated by Crank (1979), is:
@@
For slow loading, the analytical solution for advection only ($\mathrm{Pe}^{-1}=\alpha=0$) is
$c({x},t)=\frac{1}{2}\bigg{\{}\tanh{\bigg{[}{s}\left(({x}-1)\frac{\phi_{f,0}}{\phi_{f}}+{l}\right)}\bigg{]}+1\bigg{\}}.$ (51)
@@ Appendix C Dispersive flux
In this Appendix, we justify the choice of fixing $\mathrm{Pe}^{-1}$ and $\alpha$ to the specific baseline values reported in table LABEL:tab:quantitative. We do so by quantifying the strength of dispersion relative to diffusion using the ratio of the dispersive solute flux $q_{\mathrm{disp}}$ to the diffusive solute flux $q_{\mathrm{diff}}$ which we define as
@@
Figure 17: (a) Smallest and (b) largest values of $q_{\mathrm{disp}}/q_{\mathrm{diff}}$ considered in this study. The flux is plot at ten different values of $X=x-u_{s}(X,t)$ from 00 to $1$ (dark to light) during one cycle. We distinguish between the loading half of the cycle ($\dot{a}>0$; solid curves) and the unloading half of the cycle ($\dot{a}<0$; dashed curves).
@@ Appendix D Initial porosity and solute strip width
Decreasing $\phi_{f,0}$ leads to larger oscillations in fluid flux for a given solid velocity (see (22)) In this way, both advection and hydrodynamic dispersion are enhanced. However, the resulting impact on travel distance in figure 18 is relatively small because dispersion is relatively weak for these parameters.
Increasing the initial solute amount leads to a sharp increase in the travel distance: this is generally expected because the total concentration of solute is higher and therefore all the transport mechanisms are amplified. In addition, the fluid flux increases monotonically in magnitude from right to left, so larger values of $l$ expose the solute front to stronger advection and dispersion. However, note that a wider solute strip is associated with a lower degree of mixing because the variance of the solute compared to the initial variance is lower when there is more solute (see equation 38).